Prove by using distance formula that the points P(1,2,3) , Q(-1,-1,-1) and R(3,5,7) are collinear.

Grade:11

1 Answers

g v subba reddy
7 years ago
distance between pand q =29½ =5.385
distance between q and r=1161/2=10.77
distance between pand r=291/2=5.385
therefore here pq+pr=qr
so they are collinear